OpenStudy (anonymous):

-2x-5y+z=15 4x-y+3z=15 3x-2y+2z=13 Use the substitution method to solve the system what would be the equations left if z was isolated first? Please help I'm so lost

OpenStudy (anonymous):

Solve the first equation in terms of z and then put that value of z into the second 2 equations. Simplify, and now you have 2 equations in 2 unknowns.

OpenStudy (anonymous):

I got 10x-16y=-30 7x-12y=-15 I don't think its right though

OpenStudy (anonymous):

-2x-5y+z=15 eq1) mult by 3-> -6x-15y+3z=45 4x-y+3z=15 eq2) - (4x - y +3z=15) ------------------ -10x -14y = 30 eq 4) -2x-5y+z=15 eq1) mult by 2 -> -4x -10y+2z=30 3x-2y+2z=13 eq3) -> -(3x -2y +2z=13) ------------------ -7x -8y = 17 eq 5) eq 4) mult by 7 --> -70x -98y = 210 eq 5) mult by 10 -> -(-70x -80y=170) ------------------ -18y=40 y=-40/18 y=-20/9 ....ans sub y into eq 4) --> -10x -14(-20/9) =30 x=1/9 .....ans sub x and y into eq1) --> -2(1/9) -5(-20/9) +z=15 gives z=4.111 .......ans

OpenStudy (anonymous):

good luck and have fun ...lol